OpenStudy (anonymous):

This is a quadratic equation and I need to make a graph. I think have the right answer but I need to have a perfect shape "U" and something is off. Please help y=x^2-4(x). x=-1, x=0, x=1, x=2 x=3 x=4 x=5

OpenStudy (anonymous):

A plot is attached.

OpenStudy (anonymous):

Thanks!!

OpenStudy (anonymous):

y = x^2-4(x). for x=-1, y = (-1)^2 - 4(-1) = 1 + 4 = 5 so pair (x,y) = (-1, 5) for x=0, y = (0)^2 -4(0) = 0 so pair (x,y) = (0, 0) for x=1, y = (1)^2 -4(1) = 1 - 4 = -3 so pair (x,y) = (1, -3) for x=2, y = (2)^2 -4(2) = 4 - 8 = -4 so pair (x,y) = (2, -4) for x=3, y = (3)^2 -4(3) = 9-12 = -3 so pair (x,y) = (3, -3) for x=4, y = (4)^2 -4(4) = 16-16 = 0 so pair (x,y) = (4, 0) for x=5, y = (5)^2 -4(5) = 25-20 = 5 so pair (x,y) = (5, 5)

OpenStudy (anonymous):

now plot the points and you will get yr graph.......

OpenStudy (anonymous):

Thanks!! I had everything but (1-, 5) Wow I just realized that I was subtracting for y = (-1)^2 - 4(-1) = 1 + 4 = 5 instead of adding. Thanks again.

OpenStudy (anonymous):

U r welcome...☺